How many revolutions does the drive shaft of a truck engine make during its lifetime? Inputs: Let's suppose that the truck travels a million km (1 billion m) at an average speed of 80 km per hour. The engine turns at a rate of 2000 revolutions per minute. Hint: use the 'multiply by 1' technique and be careful: it's not difficult. However, there are several factors involved, so we suggest that you write your answer down formally, showing all units in the working, the way we did in the lesson.

Give your answer as a number (not a power of ten). For example, for 40 million, write 40000000. Remember to use the appropriate number of significant figures.
1
Expert's answer
2020-04-06T09:19:13-0400

As per the question,

Distance travel by the truck ="10^9 m"

Average speed (v)"=80 km/hr=80\times \dfrac{1000}{3600}=\dfrac{200}{9}m /sec"


turn rate "=2000 revolution/ min=\dfrac{2000}{60} revolution/sec"

time "t=\dfrac{10^9\times 9}{200} sec"

Hence total number of revolution =time taken in one revolution x total time

"=\dfrac{2000}{60}\times \dfrac{10^9\times 9}{200}"

"=15\times 10^8"